OpenStudy (a1234):

Amanda wants to shorten the wheelchair ramp at her front door. The current ramp is 13 feet along the horizontal and the step is 9 inches high. If she doubles the measure of the angle that the ramp makes with the walkway, how long will the new ramp be to the nearest inch?

OpenStudy (a1234):

What to do for "If she doubles the measure of the angle that the ramp makes with the walkway, how long will the new ramp be to the nearest inch?"

OpenStudy (perl):



OpenStudy (a1234):

So what would doubling the measure of the angle be?

OpenStudy (perl):

$$\sin \theta = \frac{9}{156} \\~\\\theta = \arcsin \left( \frac{9}{156} \right) \\~\\ 2\theta = 2\cdot \arcsin \left( \frac{9}{156} \right) $$

OpenStudy (perl):



OpenStudy (a1234):

Is there a way to do it without arcsin?

OpenStudy (a1234):

We have not used it before

TheSmartOne (thesmartone):

arcsin = sin^-1